The random variable Y = LogX has N(10,4) distribution. Find (a) The pdf of X (b) Mean and variance of X (c) P ( X ≤ 1000




1
Expert's answer
2022-04-20T02:34:32-0400

a). The pdf of "X" is a density of "X". The variable "Y=log\,X" has a normal distribution with parameters "10" and "4". Variable "X>0", since it appears under the function "log". Therefore, "P(X\leq0)=0". The distribution of "X" is similar to the log-normal distribution.

We have: "P(log\,X\leq z)=P(Y\leq z)=\int_{-\infty}^zg(y)dy," where "g(y)=\frac{1}{2\sqrt{2\pi}}e^{-\frac12\left(\frac{y-10}{2}\right)^2}". "P(X\leq s)=P(log\,X\leq log\,s)=\int_{0}^{log(s)}g(y)dy". The derivative(due to the Leibniz’s Theorem) is: "f(s)=(P(X\leq s))'=g(log(s))\frac{1}{s\,\,ln(10)}=\frac{1}{(2\sqrt{2\pi})\,\,(ln(10))s}e^{-\frac12\left(\frac{log(s)-10}{2}\right)^2}".

Thus, the pdf of "X" is: "f(s)=\frac{1}{(2\sqrt{2\pi})\,\,(ln(10))s}e^{-\frac12\left(\frac{log(s)-10}{2}\right)^2}"

b). "E[X]=\int_{0}^{+\infty}\frac{1}{(2\sqrt{2\pi})\,\,(ln(10))}e^{-\frac12\left(\frac{log(s)-10}{2}\right)^2}ds." Make the change of variables: "x=ln(s)". Then, we receive: "s=e^x". "E[X]=\int_{-\infty}^{+\infty}\frac{e^x}{(2\sqrt{2\pi})\,\,(ln(10))}e^{-\frac12\left(\frac{\frac{x}{ln(10)}-10}{2}\right)^2}dx=e^{\frac1{8}(a^2-100)}\int_{-\infty}^{+\infty}\frac{1}{2\,ln\,(10)\sqrt{2\pi}\,\,}e^{-\frac12\left({\frac{x-a\,ln(10)}{2\,ln(10)}}\right)^2}dx,"

where "a=4\,ln(10)+10". The last integral is equal to "1" (it becomes the Gaussian integral after simple changes). Thus, we get: "E[X]=e^{\frac1{2}((2\,ln(10)+5)^2-25)}".

"Var[X]=E[X^2]-(E[X])^2".

"E[X^2]=\int_{0}^{+\infty}\frac{s}{(2\sqrt{2\pi})\,\,(ln(10))}e^{-\frac12\left(\frac{log(s)-10}{2}\right)^2}ds." Make the change "s=e^x". We get: "E[X^2]=\int_{-\infty}^{+\infty}\frac{e^{2x}}{(2\sqrt{2\pi})\,\,(ln(10))}e^{-\frac12\left(\frac{\frac{x}{ln(10)}-10}{2}\right)^2}dx=e^{\frac1{8}(b^2-100)}\int_{-\infty}^{+\infty}\frac{1}{2\,ln\,(10)\sqrt{2\pi}\,\,}e^{-\frac12\left({\frac{x-b\,ln(10)}{2,ln(10)}}\right)^2}dx,"

where "b=8\,ln(10)+10." Thus, "Var[X]=e^{\frac1{2}((4ln(10)+5^2)-25)}-e^{((2\,ln(10)+5)^2-25)}".

We received: "E[X]=e^{\frac1{2}((2\,ln(10)+5)^2-25)}", "Var[X]=e^{\frac1{2}((4ln(10)+5^2)-25)}-e^{((2\,ln(10)+5)^2-25)}"

c). "P(X\leq1000)=\int_{0}^{1000}\frac{1}{(2\sqrt{2\pi})\,\,(ln(10))s}e^{-\frac12\left(\frac{log(s)-10}{2}\right)^2}ds\approx0.0002" (it is rounded to 4 decimal places)